Math

(1 point) Find the volume of the solid obtained by rotating the region in the first quadrant bounded by \(y=x^{7}, y=1\), and the \(y\) -axis around the \(y\) -axis.

Volume =
Answer #1

The statement of disk method/ washer method when region is revolved about Y axis is:

Disk method / washer was her method / Revolution about Y anie ) X are integrable Let flufa : [a,b] > R functi off Let region

Now let's begin to solve given question as:

in first (0) Find the volume of solid olid when region quadrant in bounded by y=x y=x7, y=1 and Yaxin about Η ακτλ. Soln! letand Region in bounded between y=0 f g=1 OLYEI } want to we DiAK method, we select on yecoil since we व faly)= y fily) = 0 2 -2+7 C → 7 I cs I. 247 7 ت له whace FIT y [*]. P.14 V = Volume 7T 11-0] - l volume FIT 9 Anwen or volume 2.44 35

I am attaching the graphs drawn in desmos for better understanading

III (0.1) (1.1) -10 -5 o 5 10 (0,0) -5 + Х y=x? Х X=0 -10 10 3 y=1 -10 10 4

Note: for any query or explanataion in any step, kindly mention in comment section. I will assist you ASAP.